OpenStudy (anonymous):

What angle does the line P(7,-4) and Q(1,6) make with the positive x-axis?

OpenStudy (anonymous):

aractan(10/-6)

OpenStudy (amistre64):

(7,-4) -(1, 6) ------- <6,10> = 2<3,5>

OpenStudy (amistre64):

<3,5> <1,0> ------ 3+0 = 3 \[cos(a)=\frac{3}{\sqrt{34}}\] \[cos^{-1}(\frac{3}{\sqrt{34}})\ :)\]

OpenStudy (amistre64):

just to be different lol

OpenStudy (anonymous):

The answer in the back of the book says it's 122

OpenStudy (amistre64):

i dropped a negative, doh!

OpenStudy (amistre64):

it should be about 59 degrees or so

OpenStudy (anonymous):

it is in 2nd quardant take arc tan(10/6) then 180-arc tan(10/6)

OpenStudy (amistre64):

thats better

OpenStudy (anonymous):

Since it's in the 2nd quadant you have to subtract 180?